OpenStudy (anonymous):

If the spinner shown is spun twice, what is the probability that it will land on 1 both times?

OpenStudy (anonymous):

how many sections?

OpenStudy (anonymous):

will you describe the spinner at least?

OpenStudy (llort):

1/n^2

OpenStudy (amistre64):

1/n * 1/n = 1/n^2

OpenStudy (anonymous):

8

OpenStudy (amistre64):

\begin{array}l\color{red}{\text{1}}\color{orange}{\text{/}}\color{#9c9a2e}{\text{6}}\color{green}{\text{4}}\color{blue}{\text{ }}\color{purple}{\text{t}}\color{purple}{\text{h}}\color{red}{\text{e}}\color{orange}{\text{n}}\color{#9c9a2e}{\text{ }}\color{green}{\text{:}}\color{blue}{\text{)}}\color{purple}{\text{}}\end{array}

OpenStudy (llort):

1/8^2=1/64

OpenStudy (llort):

ninja'd